# A Semi-Supervised Learning Method for the Identification of Bad Exposures in Large Imaging Surveys

### v1

@@ Abstract
As the data volume of astronomical imaging surveys rapidly increases, traditional methods for image anomaly detection, such as visual inspection by human experts, are becoming impractical. We introduce a machine-learning-based approach to detect poor-quality exposures in large imaging surveys, with a focus on the DECam Legacy Survey (DECaLS) in regions of low extinction (i.e., $E(B-V)<0.04$). Our semi-supervised pipeline integrates a vision transformer (ViT), trained via self-supervised learning (SSL), with a k-Nearest Neighbor (kNN) classifier. We train and validate our pipeline using a small set of labeled exposures observed by surveys with the Dark Energy Camera (DECam). A clustering-space analysis of where our pipeline places images labeled in “good” and “bad” categories suggests that our approach can efficiently and accurately determine the quality of exposures. Applied to new imaging being reduced for DECaLS Data Release 11, our pipeline identifies 780 problematic exposures, which we subsequently verify through visual inspection. Being highly efficient and adaptable, our method offers a scalable solution for quality control in other large imaging surveys.
@@ 1 Introduction
The Astro2020 decadal survey (National Academies of Sciences, Engineering, and Medicine, 2023) highlighted the increasing importance of advanced pipelines for processing data from large astronomical surveys, and how scientific results will crucially depend on these pipelines. Identifying bad exposures is a key step in any pipeline to optimize the quality of images from sky surveys. However, this step can often be laborious. Traditionally, the bad exposures can be labeled in two ways. First, the observers at the telescope can record the exposure condition and quality during observation. Some issues, such as shaking due to strong winds, can be noted. Further, some common issues, such as hot pixels and cosmic rays, can be detected and fixed using algorithms and available software (van Dokkum, 2001; Astropy Collaboration et al., 2022). Some other issues, like seeing and Point Spread Function (PSF) variations, are usually detectable during photometric reductions (Lang et al., 2016). Beyond this, more specific problems, such as ghosting and scattered light, have been historically difficult to handle and have often required special algorithmic treatment. Although some solutions to these problems exist (Patterson & Wells, 2003; Kent, 2013), they are usually built for a specific telescope or camera. In this case, astronomers have to visually inspect exposures to determine which might be problematic and/or evaluate the image quality during the overall data reduction process. Astronomers often also have to explicitly construct separate models or metrics to study each problematic imaging systematic.
@@
As larger surveys proceed to everyday operations, the number of exposures in their associated imaging campaigns is likely to become too high for a human expert to visually inspect every image. Further, as the number of pixels achieved by cameras increases, features in images are likely to become more complex, which will make the identification of bad exposures even more taxing. As an example, the LSSTCam on the Vera C. Rubin telescope utilizes a 3.2-gigapixel CCD detector and can observe 10,000 square degrees of sky in a single filter every three clear nights (Ivezić et al., 2019). Despite the increase in scope and complexity of future imaging surveys, the quality of exposures in large campaigns will remain essential to downstream tasks, such as astrometric (e.g., Lang et al., 2009), and photometric (e.g., Stubbs & Tonry, 2006) measurements, and target selection to facilitate follow-up studies (e.g., Myers et al., 2023). The bottom line is that unintentionally including bad exposures in the datasets produced by imaging surveys will always be detrimental to downstream scientific results. Therefore, understanding and identifying problematic exposures will be a difficult but crucial element of next-generation imaging surveys.
@@
While large astronomical surveys have been expanding in volume and quality, more efficient data processing and analysis algorithms have already started to become desirable to take the burden off human inspectors. For example, Zhang & Bloom (2020) and Chang et al. (2021) used Convolutional Neural Networks (CNNs) to, respectively, identify cosmic rays in Hubble Space Telescope images, and detect ghosts and scattered light in Dark Energy Survey (DES) images; Tanoglidis et al. (2021) applied a method based on a Mask Region-Based Convolutional Neural Network (R-CNN) to predict the location of ghosts and scattered light in Dark Energy Camera (henceforth DECam; Flaugher et al., 2015) exposures; and Parisot et al. (2023) applied existing image quality assessment (IQA) algorithms developed for general computer vision applications to score images drawn from the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Surveys (dubbed LS in the following text; Dey et al., 2019), based on a predefined set of bad images.
@@
In this work, we focus on using recently developed machine learning (henceforth ML) algorithms to identify bad exposures from DECam. This is a complex problem as each DECam exposure comprises 62 individual CCDs (or 61 CCDs with N30 masked out [^1] More details can be found at: https://noirlab.edu/science/programs/ctio/instruments/Dark-Energy-Camera/Status-DECam-CCDs), and problems can be present both on large scales (i.e., the entire exposure) and small scales (i.e., individual CCDs, or even groups of pixels). Some issues with the DECam exposures have been mitigated by the DECam Community Pipeline (CP; Valdes et al., 2014), which is a data reduction pipeline that processes DECam images and produces calibrated images and catalogs. However, some issues, such as ghosting and scattered light, are not fully addressed by the CP, and thus require further investigation. To address this complexity, our pipeline adopts an ML approach based on a Vision Transformer (ViT; Dosovitskiy et al., 2021). ViT is a deep learning model based on the transformer architecture (Vaswani et al., 2017) to capture complex patterns in images, often outperforming traditional convolutional neural networks (CNNs) when trained on large datasets. In particular, we use a pretrained existing ViT in the context of version 2 of the self-DIstillation with No Labels (DINOv2) framework (Caron et al., 2021; Oquab et al., 2023). The original ViT was trained with ImageNet (Deng et al., 2009), and we apply that trained ViT to generate representations of an image, generally referred to as a vector “embedding.” An embedding is essentially a representation of the features of an image in a high-dimensional parameter space. We then use a k-Nearest Neighbor (kNN) algorithm to classify these embeddings after dimension reduction, and assign categories to images based on a labeled training set.
@@
We aim to simultaneously detect multiple potential problems in an exposure, and use a pipeline that only requires a relatively small training and testing dataset, typically comprising less than 5% of the whole exposure dataset. As a purely data-driven method, our model can be adapted to other telescopes, although we choose to train our model using a dataset from DECam. Our pipeline is also relatively fast for inference. To tackle the large volume of data, our framework uses Graphics Processing Units (GPU). GPUs are a highly adaptable and efficient hardware for big datasets and are widely used in ML fields (Steinkraus et al., 2005). Moreover, we take advantage of multi-GPU parallel processing to achieve higher inference speed as long as the model can be fitted into each single GPU.
@@ 2 Dataset
The exposures in our dataset are mostly drawn from the LS Data Releases 9, 10, and 11 which we henceforth refer to as DR9, DR10 and DR11. The LS comprise three surveys: the Beijing-Arizona Sky Survey (BASS; Zou et al., 2017), Mayall z-band Legacy Survey (MzLS), and DECam Legacy Survey (DECaLS; Dey et al., 2019). [^2] We use LS to refer specifically to the DECaLS portion of the LS in the following text. We focus our analysis on DECam images from DECaLS collected with the Victor M. Blanco 4-meter Telescope [^3] https://noirlab.edu/public/programs/ctio/victor-blanco-4m-telescope/ at the Cerro Tololo Inter-American Observatory (CTIO) in Chile. DR11 also includes exposures from the DECam Local Volume Exploration Survey (DELVE; Drlica-Wagner et al., 2021, 2022), which are incorporated in our dataset. All of the images we consider in this paper are inputs to, or have been reduced by, the DECam Community Pipeline (CP; Valdes et al., 2014, see Section 2.4). A small fraction of the exposures are labeled by experts as problematic across a range of different categories. These labels come from three major sources: the DECam bad exposure list (see Section 2.1), previous experts’ visual inspection labels from DELVE, and a wider list of inspections flagged for the DES (Section 2.2 and Appendix A). In addition, we use bad exposure flags developed for DR9 and DR10 to narrow down the selection of good exposures (see Section 2.3). The numbers of categories that can be used as labels, and the numbers of bad exposures provided by each source are listed in Table 1. There are 256,701 exposures in DR9 and DR10, and 18,315 exposures are labeled with at least 1 CCD image as bad by human experts, which is about 7% of the total exposures.
Exposures from DR9 and DR10 are used in our training and testing process. New exposures that will be added for DR11, as well as old exposures from DR9 and DR10 that are reprocessed for DR11, are included in our inference dataset. Our task is to identify bad exposures from the unlabeled exposures based on the labeled exposures, which is a problem that is well-suited to a multi-class classification-task design. With this problem design, a balanced dataset is essential in the training and evaluation process for the classification. We therefore divide our dataset into three exclusive samples:
@@ 2.1 DECaLS Bad Exposure List
The DECaLS bad exposure list records 3,014 bad exposures identified during DECaLS observations and recorded in an observing log file named bad_expid.txt [^4] https://github.com/legacysurvey/legacypipe/blob/DR10.2/py/legacyzpts/data/decam-bad_expid.txt. The issues recorded in the log file can be divided into three major categories: exposure quality, observing conditions, and instrument issues. The log file uses three main flags, which are “possibly saturated” (denoted as 1 in bad_expid.txt), “seeing” (denoted as 2), and “other sources” (denoted as 3). We used the first two flags directly to indicate bad exposures. These issues span a large domain and each exposure may have slightly different severity. Therefore, to clearly define the bad exposures in this list, we consider exposures in the list satisfying any of the following conditions as bad:
  • the exposure flag is set to 1 or 2
  • the exposure flag is set to 3, and either one of the conditions:
    • expfactor is included in the comment and is less than 1. expfactor means the exposure factor of an exposure, and it is used to indicate the amount of light
    • the comment contains one of the keywords: telescope moving, tracking, PSF, focus.
The “other source” flag usually has associated comments, which we parsed to capture keywords consistent with the above categories. We expect that seeing, transparency, and telescope instability will influence scales across an entire exposure, whereas other categories may be limited to smaller scales that are more consistent with a single CCD.
@@ 2.2 Expert-Labeled Exposures
The expert labels we use are provided by astronomers who visually inspected the exposures and evaluated their quality, in particular in the context of DELVE and DES. The aim of DES (DES Collaboration, 2016) is to survey 5000 deg2 of the southern sky in five optical bands ($grizY$). The DES expert-labeled bad exposure list is based on the DES Year 6 (Y6) data release. There are 122,531 CCD images labeled as bad from 6,667 exposures. These exposures are categorized into 61 different types of comments. Some of the comments are associated with the same imaging patterns but using different expressions or for different reasons, such as “telescope moving” and “telescope tracking failure.” We therefore combined some cases. Ultimately, we picked the top 10 most frequent issues as our bad exposure labels.
DELVE uses DECam to study ultra-faint galaxies around the Milky Way, Magellanic Clouds, and isolated dwarf galaxies in the Local Volume (Drlica-Wagner et al., 2021). The expert-labeled bad exposures we use in this paper were assembled on 2023 July 25 as an input into DELVE DR3 (Tan et al., 2025). The expert-labeled DELVE bad exposure dataset contains 51,445 problematic CCD images from 8,634 exposures. There are 37 bad categories associated with these exposures, and we picked the top 10 most frequent issues, similar to the process we used for the DES dataset.
The bad categories in DES do not perfectly match with the categories in DELVE, so we created a unified list of bad categories to combine identified CCD images from the two lists. The final number of images is shown in Table 2. These two datasets provide a large number of labeled bad exposures, which is an essential prerequisite for effectively training the classifier we use in this paper.
Throughout the rest of this paper, we will refer to the Expert-Labeled Exposures List as “the human expert list” or the list “labeled by human experts.”
@@
legacypipe555code available at https://github.com/legacysurvey/legacypipe is a pipeline based on the Tractor (Lang et al., 2016) that is used to process the LS imaging data. As part of the reduction process, legacypipe provides a set of flags to indicate the quality of each exposure (see, e.g. Dey et al., 2019, for more information). The quality flags are stored in the survey CCD FITS file as bitmasks ranging from 0 to 18, corresponding to 19 different potential issues.666e.g., https://www.legacysurvey.org/dr10/bitmasks/#ccd-cuts
We used the legacypipe flags to remove exposures in the good exposure candidate sets that are not in the DECaLS list nor labeled by human experts. This rejection process will minimize the contamination of the good exposure set by bad exposures, hopefully producing a more accurate training process.
@@ 2.4 Community Pipeline
The DECam Community Pipeline (CP; Valdes et al., 2014) is an imaging reduction pipeline that is used to process the DECaLS. The DECaLS images are processed by various CP versions, each of which targets some specific problem and, generally, a later CP version is more comprehensive and thus better than a previous one. The range of CP versions used in the LS is from V4.10 through V5.5LS, with all processing performed later than 2019-06-06.777Information on CP versions: https://desi.lbl.gov/trac/wiki/DecamLegacy/CPVersions For each exposure, the CP produces many reduced data files and diagnostics. We chose to use the ooi images [^8] For a more detailed description see https://noirlab.edu/science/index.php/data-services/data-reduction-software/csdc-mso-pipelines/pl206#tab10 produced by the CP for our training process because those images are used for stacking and generating the bricks used for LS processing and better match the large-scale background — they are therefore more representative of the eventual images used for scientific analyses. Some exposures do not have associated ooi files because the CP processing ended prematurely due to problems in the exposures. Since those exposure issues are usually very obvious, we ignored them during our training as they did not provide additional useful information.

### v2

@@ Abstract
As the data volume of astronomical imaging surveys rapidly increases, traditional methods for image anomaly detection, such as visual inspection by human experts, are becoming impractical. We introduce a machine-learning-based approach to detect poor-quality exposures in large imaging surveys, with a focus on the DECam Legacy Survey (DECaLS) in regions of low extinction (i.e., E(BV)<0.04E(B-V)<0.04). Our semi-supervised pipeline integrates a vision transformer (ViT), trained via self-supervised learning (SSL), with a k-Nearest Neighbor (kNN) classifier. We train and validate our pipeline using a small set of labeled exposures observed by surveys with the Dark Energy Camera (DECam). A clustering-space analysis of where our pipeline places images labeled in “good” and “bad” categories suggests that our approach can efficiently and accurately determine the quality of exposures. Applied to new imaging being reduced for DECaLS Data Release 11, our pipeline identifies 780 problematic exposures, which we subsequently verify through visual inspection. Being highly efficient and adaptable, our method offers a scalable solution for quality control in other large imaging surveys.
@@ 1 Introduction
The Astro2020 decadal survey (National Academies of Sciences, Engineering, and Medicine, 2023) highlighted the increasing importance of advanced pipelines for processing data from large astronomical surveys, and how scientific results will crucially depend on these pipelines. Identifying bad exposures is a key step in any pipeline to optimize the quality of images from sky surveys. However, this step can often be laborious. Traditionally, the bad exposures can be labeled in two ways. First, the observers at the telescope can record the exposure condition and quality during observation. Some issues, such as shaking due to strong winds, can be noted. Further, some common issues, such as hot pixels and cosmic rays, can be detected and fixed using algorithms and available software (van Dokkum, 2001; Astropy Collaboration et al., 2022). Some other issues, like seeing and Point Spread Function (PSF) variations, are usually detectable during photometric reductions (Lang et al., 2016). Beyond this, more specific problems, such as ghosting and scattered light, have been historically difficult to handle and have often required special algorithmic treatment. Although some solutions to these problems exist (Patterson and Wells, 2003; Kent, ), they are usually built for a specific telescope or camera. In this case, astronomers have to visually inspect exposures to determine which might be problematic and/or evaluate the image quality during the overall data reduction process. Astronomers often also have to explicitly construct separate models or metrics to study each problematic imaging systematic.
@@
As larger surveys proceed to everyday operations, the number of exposures in their associated imaging campaigns is likely to become too high for a human expert to visually inspect every image. Further, as the number of pixels achieved by cameras increases, features in images are likely to become more complex, which will make the identification of bad exposures even more taxing. As an example, the LSSTCam on the Vera C. Rubin telescope utilizes a 3.2-gigapixel CCD detector and can observe 10,000 square degrees of sky in a single filter every three clear nights (Ivezić et al., 2019). Despite the increase in scope and complexity of future imaging surveys, the quality of exposures in large campaigns will remain essential to downstream tasks, such as astrometric (e.g., Lang et al., 2009), and photometric (e.g., 46) measurements, and target selection to facilitate follow-up studies (e.g., Myers et al., 2023). The bottom line is that unintentionally including bad exposures in the datasets produced by imaging surveys will always be detrimental to downstream scientific results. Therefore, understanding and identifying problematic exposures will be a difficult but crucial element of next-generation imaging surveys.
@@
While large astronomical surveys have been expanding in volume and quality, more efficient data processing and analysis algorithms have already started to become desirable to take the burden off human inspectors. For example, Zhang and Bloom (2020) and Chang et al. (2021) used Convolutional Neural Networks (CNNs) to, respectively, identify cosmic rays in Hubble Space Telescope images, and detect ghosts and scattered light in Dark Energy Survey (DES) images; Tanoglidis et al. (2021) applied a method based on a Mask Region-Based Convolutional Neural Network (R-CNN) to predict the location of ghosts and scattered light in Dark Energy Camera (henceforth DECam; Flaugher et al., 2015) exposures; and Parisot et al. (2023) applied existing image quality assessment (IQA) algorithms developed for general computer vision applications to score images drawn from the Dark Energy Spectroscopic Instrument (DESI) Legacy Imaging Surveys (dubbed LS in the following text; Dey et al., 2019), based on a predefined set of bad images.
@@
In this work, we focus on using recently developed machine learning (henceforth ML) algorithms to identify bad exposures from DECam. This is a complex problem as each DECam exposure comprises 62 individual CCDs (or 61 CCDs with N30 masked out [^1] More details can be found at: https://noirlab.edu/science/programs/ctio/instruments/Dark-Energy-Camera/Status-DECam-CCDs), and problems can be present both on large scales (i.e., the entire exposure) and small scales (i.e., individual CCDs, or even groups of pixels). Some issues with the DECam exposures have been mitigated by the DECam Community Pipeline (CP; Valdes et al., 2014), which is a data reduction pipeline that processes DECam images and produces calibrated images and catalogs. However, some issues, such as ghosting and scattered light, are not fully addressed by the CP, and thus require further investigation. To address this complexity, our pipeline adopts an ML approach based on a Vision Transformer (ViT; Dosovitskiy et al., 2021). ViT is a deep neural network that splits an image into small patches and learns patterns from them. It is based on the transformer architecture (Vaswani et al., 2017) to capture complex patterns in images, often outperforming traditional convolutional neural networks (CNNs) when trained on large datasets. In particular, we use a pretrained existing ViT in the context of version 2 of the self-DIstillation with No Labels (DINOv2) framework (Caron et al., 2021; Oquab et al., 2023), a self-supervised method that learns from images without human-provided labels. The original ViT was trained with ImageNet (Deng et al., 2009), and we apply that trained ViT to generate a numeric vector “embedding” that summarizes the image’s content. An embedding is essentially a representation of the features of an image in a high-dimensional parameter space. We then use a k-Nearest Neighbor (kNN) algorithm to classify these embeddings after dimension reduction, and assign categories to images based on a labeled training set.
@@
We aim to simultaneously detect multiple potential problems in an exposure, and use a pipeline that only requires a relatively small training and testing dataset, typically comprising less than 5% of the whole exposure dataset. Because our model learns directly from the data, it can be adapted to other telescopes, although we choose to train our model using a dataset from DECam. Our pipeline is also relatively fast for inference. To tackle the large volume of data, our framework uses Graphics Processing Units (GPU). GPUs are a highly adaptable and efficient hardware for big datasets and are widely used in ML fields (Steinkraus et al., 2005). Moreover, we take advantage of multi-GPU parallel processing to achieve higher inference speed as long as the model can be fitted into each single GPU.
@@ 2 Dataset
The exposures in our dataset are mostly drawn from the LS Data Releases 9, 10, and 11 which we henceforth refer to as DR9, DR10 and DR11. The LS comprise three surveys: the Beijing-Arizona Sky Survey (BASS; Zou et al., 2017), Mayall z-band Legacy Survey (MzLS), and DECam Legacy Survey (DECaLS; Dey et al., 2019). [^2] We use LS to refer specifically to the DECaLS portion of the LS in the following text. We focus our analysis on DECam images from DECaLS collected with the Victor M. Blanco 4-meter Telescope [^3] https://noirlab.edu/public/programs/ctio/victor-blanco-4m-telescope/ at the Cerro Tololo Inter-American Observatory (CTIO) in Chile. DR11 also includes exposures from the DECam Local Volume Exploration Survey (DELVE; Drlica-Wagner et al., 2021, 2022), which are incorporated in our dataset. All of the images we consider in this paper are inputs to, or have been reduced by, the DECam Community Pipeline (CP; Valdes et al., 2014, see Section 2.4). A small fraction of the exposures are labeled by experts as problematic across a range of different categories. These labels come from three major sources: the DECam bad exposure list (see Section 2.1), previous experts’ visual inspection labels from DELVE, and a wider list of inspections flagged for the DES (Section 2.2 and Appendix C). In addition, we use bad exposure flags developed for DR9 and DR10 to narrow down the selection of good exposures (see Section 2.3). The numbers of categories that can be used as labels, and the numbers of bad exposures provided by each source are listed in Table 1. There are 256,701 exposures in DR9 and DR10, and 18,315 exposures are labeled with at least 1 CCD image as bad by human experts, which is about 7% of the total exposures.
Our task is to identify bad exposures from the unlabeled exposures based on the labeled exposures, which is a problem that is well-suited to a multi-class classification-task design. With this problem design, a balanced dataset is essential in the training and evaluation process for the classification. We therefore divide our dataset into three exclusive samples:
@@ 2.1 DECaLS Bad Exposure List
The DECaLS bad exposure list records 3,014 bad exposures identified during DECaLS observations and recorded in an observing log file named bad_expid.txt [^4] https://github.com/legacysurvey/legacypipe/blob/DR10.2/py/legacyzpts/data/decam-bad_expid.txt. The issues recorded in the log file can be divided into three major categories: exposure quality, observing conditions, and instrument issues. The log file uses three main flags, which are “possibly saturated” (denoted as 1 in bad_expid.txt), “seeing” (denoted as 2), and “other sources” (denoted as 3). As “other sources” denotes a wide range of issues of varying severity, we consider exposures in the list satisfying the following criteria as “bad”:
  • the exposure flag is set to 1 or 2; or
  • the exposure flag is set to 3, and either:
    • expfactor is included in the comment and is less than 1; or
    • the “other sources” comment contains one of the keywords: telescope moving, tracking, PSF or focus.
Here, expfactor is an exposure factor defined as the ratio between the effective exposure time and the nominal exposure time. It is used to indicate the amount of light collected by the CCDs during the exposure after accounting for the transparency, seeing and instrument throughout. A value lower than 1 indicates that the exposure is not optimal.
@@ 2.2 Expert-Labeled Exposures
The expert labels we use are provided by astronomers who visually inspected the exposures and evaluated their quality, in particular in the context of DELVE and DES. The aim of DES (DES Collaboration, 2016) is to survey 5000 deg2 of the southern sky in five optical bands ($grizY$). The DES expert-labeled bad exposure list is based on the DES Year 6 (Y6) data release. There are 122,531 CCD images labeled as bad from 6,667 exposures. These exposures are categorized into 61 different types of comments, but we combined comments that signified similar issues. For example, we incorporated both “HeavyClouds” and “Clouds and transparency” into the Clouds_Transparency category.
DELVE uses DECam to study ultra-faint galaxies around the Milky Way, Magellanic Clouds, and isolated dwarf galaxies in the Local Volume (Drlica-Wagner et al., 2021). The expert-labeled bad exposures we use in this paper were assembled on 2023 July 25 as an input into DELVE DR3 (Tan et al., 2025). The expert-labeled DELVE bad exposure dataset contains 51,445 problematic CCD images from 8,634 exposures. There are 37 bad categories associated with these exposures.
The bad categories in DES do not perfectly match those in DELVE, so we created a unified list of bad categories to combine identified CCD images from the two surveys. After combining similar cases, we chose the 10 most frequent issues appearing in the DES and/or DELVE datasets as our bad exposure labels. The final number of images is shown in Table 4. These two datasets provide a large number of labeled bad exposures, which is an essential prerequisite for effectively training the classifier we use in this paper. Throughout the rest of this paper, we will refer to the Expert-Labeled Exposures List as “the human expert list”
@@
legacypipe555code available at https://github.com/legacysurvey/legacypipe is a pipeline based on the Tractor (Lang et al., 2016) that is used to process the LS imaging data. As part of the reduction process, legacypipe provides a set of flags to indicate the quality of each exposure (see, e.g. Dey et al., 2019, for more information). The quality flags are stored in the survey CCD FITS file as integers that combine bits ranging from $2^{0}$ to $2^{18}$ where each individual bit indicates a specific condition.666e.g., https://www.legacysurvey.org/dr10/bitmasks/#ccd-cuts
We used the legacypipe flags to further prune exposures that are not in the DECaLS list or the human expert list. The total number of exposures with bad legacypipe flags is 33.21% of the total exposures in DR9 and DR10. 46.54% of all exposures include at least one CCD with a bad legacypipe flag. Applying the legacypipe flag cut removes these exposures and minimizes the contamination of the good exposure set by bad CCDs. We only selected exposures with no bad legacypipe flags as good exposures in our training and testing datasets. This will improve the purity of the good exposure set, leading to a more robust training and testing process.
@@ 2.4 Community Pipeline
The DECam Community Pipeline (CP; Valdes et al., 2014) is an imaging reduction pipeline that is used to process the DECaLS. The range of CP versions used in the LS is from V4.10 through V5.5LS, with all processing performed later than 2019-06-06.777Information on CP versions: https://desi.lbl.gov/trac/wiki/DecamLegacy/CPVersions For each exposure, the CP produces many reduced data files and diagnostics. We chose to use the ooi images [^8] For a more detailed description see https://noirlab.edu/science/index.php/data-services/data-reduction-software/csdc-mso-pipelines/pl206#tab10 produced by the CP for our training process because those images are used for stacking and generating the small regions of the sky (“bricks”) that are defined to facilitate LS processing. These ooi images are better calibrated to the large-scale background than other LS products, and best represent the final LS science images. Some exposures do not have associated ooi files because the CP processing ended prematurely due to problems in the exposures. Since those exposure issues are usually very obvious, we ignored them during our training as they did not provide additional useful information.